Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB SQS SXP XO Migrated2Zulip Archived2GitlabPages Archived2PDF not2MigrateOrArchive MigratedByGroup
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip DET Sample Environment SPB_mig
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 108     Entry time: 06 Apr 2020, 09:50
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: notes Dipatch Meeting April 6th, 2020 
CTRL:
- It was remotely observed that some cameras were consuming too much memory. By remotely re-starting these devices the issue was solved.  
- No calls to the OCD during the weekend.

EEE:
- No issues to report;
- Planning a meeting with Instruments/CTRL/DA to discuss long term tasks that have been working on during this reduced operation period.

VAC:
- Nothing to report; All running in good state.

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM:
- Nothing to report.

XRO:
- After the recent power glitch: XRO has checked the basic parameters of special motors which presented problems in the past associated to power glitches;
- At this time, the main basic parameters remain OK (note that only few devices were checked - the ones which presented problems in the past);
- During the EuXFEL restart a deeper look on all devices has to be taken together with the instruments. 

SPB:
- Nothing to report.

FXE:
- Nothing to report.

MID:
- Nothing to report.

HED:
- Nothing to report.

	 
ELOG V3.1.4-7c3fd00